# DevBench: A Comprehensive Benchmark for Software Development

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced their coding capabilities. However, existing benchmarks predominantly focused on simplified or isolated aspects of programming, such as single-file code generation or repository issue debugging, falling short of measuring the full spectrum of challenges raised by real-world programming activities. To this end, we propose DevBench, a comprehensive benchmark that evaluates LLMs across various stages of the software development lifecycle, including software design, environment setup, implementation, acceptance testing, and unit testing. DevBench features a wide range of programming languages and domains, high-quality data collection, and carefully designed and verified metrics for each task. Empirical studies show that current LLMs, including GPT-4-Turbo, fail to solve the challenges presented within DevBench. Analyses reveal that models struggle with understanding the complex structures in the repository, managing the compilation process, and grasping advanced programming concepts. Our findings offer actionable insights for the future development of LLMs toward real-world programming applications. Our benchmark is available at https://github.com/open-compass/DevBench.
Machine Learning, ICML
Figure 1: Our DevBench features multiple stages of software development, including software design, environment setup, implementation, and testing (both acceptance and unit testing).
[…]
@@
Table 1: Task design in DevBench. $\dagger$}: following our modular evaluation protocol, the input for each task are reference. $\ddagger$}: includes UML class and sequence diagrams. $\ast$}: implementation source code is optional for acceptance testing.
1 Introduction
@@
Given its practical value and reasoning challenges, programming has become an important domain to deploy and evaluate large language models (LLMs), leading to popular products like GitHub Copilot and research benchmarks like HumanEval (Chen et al., 2021) and APPS (Hendrycks et al., 2021). While these earlier coding tasks focused on generating a single code file or even a single method conditioned on simple instructions, recent works such as SWE-bench (Jimenez et al., 2023) and RepoBench (Liu et al., 2023b) propose new evaluations that challenge LLMs with repository-level coding tasks, which feature longer, more involved NL2Code problems, that capture practical challenges in the real world. Still, these benchmarks concentrate on narrow aspects of software development. To this end, datasets and metrics that comprehensively evaluate software development as a multi-phase set of tasks are notably absent.
@@
To address these shortcomings and fill this gap, we propose DevBench, a novel benchmark that mirrors real-world software development. DevBench generally evaluates models on the task of constructing a multi-file codebase starting from a product requirement document (PRD) of detailed specifications. Subscribing to the traditional Waterfall software development model (Royce, 1987), DevBench breaks down this process into a diverse set of inter-related development stages, i.e., software design, environment setup, implementation, acceptance and unit testing, as visualized in Figure 1 and Table 1. Across all tasks, DevBench is the first to evaluate models’ software design and environment setup capabilities.
It is not trivial to collect a dataset for DevBench, as open-sourced repository data often lack essential components, especially the design documents and testing programs. The DevBench dataset is made up of 22 curated repositories in 4 languages (Python, C/C++, Java, JavaScript). These codebases cover a variety of domains, including machine learning, web service, command line utilities, etc. By encompassing the multi-faceted, interconnected steps of software development under a single framework, DevBench provides a holistic view of the capabilities of LLMs for automated software production, moving beyond the conventional, singular focus on code completion.
@@
We also develop a baseline system by employing ChatDev (Qian et al., 2023a, b) as the foundational framework to benchmark state-of-the-art pretrained code models. We find that all models fail to solve the challenges presented within DevBench. GPT-4-Turbo achieves the highest scores amongst all evaluated models, yet it obtains less than 10% on our repository-level implementation task. Other tasks prove relatively more manageable for models; however, even GPT-4-Turbo struggles to attain scores exceeding 40% on any given task. For instance, models generally fail to generate executable testing code, with oracle test scores falling below 40%. Despite this, the generated testing code demonstrates potential in terms of coverage score, achieving as high as 79.4% when it is executable. Furthermore, our investigation into different prompting methods shows that only prompts supplemented with external information, execution feedback in our context, yield notable and consistent improvements. We employ LLM-as-a-Judge (Zheng et al., 2023a) for evaluating software design, where the outcomes exhibit high agreement with human annotations. Importantly, qualitative analysis shows that models demonstrate difficulties in handling Makefile and Gradle, configuring function arguments and employing advanced object-oriented programming techniques. Overall, DevBench introduces a novel challenge for existing code models, and our investigation sheds light on fundamental issues, paving the way for future research.
@@ 2 Related Work
While initial code generation datasets mainly put forth self-contained, closed form completion problems predominantly written in Python (Chen et al., 2021; Hendrycks et al., 2021; Austin et al., 2021), subsequent works have raised coding tasks’ complexity via a number of augmentations, including increasing language coverage (Zheng et al., 2023b; Cassano et al., 2023), expanding execution-based test coverage (Liu et al., 2023a; Wang et al., 2022), requiring use of dependencies (Lai et al., 2022; Ding et al., 2023; Liu et al., 2023c), proposing interactive coding environments (Yin et al., 2022; Yang et al., 2023), and constructing suites of short-form coding tasks (Lu et al., 2021; Muennighoff et al., 2023). DevBench falls most in line with recent works that introduce repository-scale coding, where a task worker must generate a code completion or patch that fixes an issue within a production-level codebase (Jimenez et al., 2023; Liu et al., 2023b; Zhang et al., 2023). While these prior works mainly focus on iterative software development, they do not investigate LLMs’ ability to create a codebase from extensive natural language project descriptions.
In automatic programming, recent works propose role-play frameworks for programming that involves communicative agents (Hong et al., 2023; Li et al., 2023; Qian et al., 2023a, b). Among them, MetaGPT (Hong et al., 2023) is the most closely related to our work, which integrates Standardized Operating Procedures into prompt sequences for streamlined workflows. Additionally, MetaGPT is characterized by structured output formats and a modular task design that resembles our work. The SoftwareDev auxiliary dataset they put forth is focused on cost, code statistics, and a general assessment of executability. In contrast, DevBench pioneers a novel evaluation approach founded on a carefully curated dataset of repositories. We define structured, thorough, and comprehensive evaluations across every phase of the software development lifecycle.
3 DevBench
In this section, we discuss the design of DevBench, including task specifications and the evaluation criteria (summarized in Table 1). Additional insights and exemplar artifacts of the tasks are presented in Appendix A. Adhering to the modular approach, our design utilizes reference inputs for each task, instead of relying on outputs generated by models from preceding stages. This strategy enables and concentrates on evaluating the efficacy of models in executing specific tasks. 111 Notably, our framework can be adeptly configured to facilitate end-to-end evaluations, utilizing the intermediate outputs generated by models across multiple tasks. Moreover, DevBench can also function in a Copilot mode, empowering human users to intervene and refine model outputs, thus enhancing the collaborative synergy between human expertise and automated systems.
@@ 3.1 Task 1: Software Design
During this phase, the model is tasked with interpreting the Product Requirements Document (PRD) to create Unified Modeling Language (UML) diagrams, utilizing the extensively recognized Mermaid syntax, alongside formulating architecture designs using hierarchical file-tree structures. The class diagram acts as a fundamental component in software system modeling, illustrating classes, their attributes, and interrelations. Sequence diagrams detail processes and objects involved and the sequence of messages exchanged to carry out the software functionality. Architecture design is aimed at crafting a structured layout of source code, compilation scripts, and auxiliary files crucial for software implementation. For exemplifications of UML diagrams and architectural frameworks, refer to Appendix A.1.
Evaluation. Since the Software Design tasks are open-ended, we employ the LLM-as-a-judge approach (Zheng et al., 2023a; Wang et al., 2023; Chiang & Lee, 2023) to conduct the automatic evaluation. The evaluation is anchored by two principal metrics: general principles and faithfulness. Detailed evaluation guidance is found in Appendix B.
@@ 3.2 Task 2: Environment Setup
In the second phase of DevBench, models are provided with the PRD, UML diagrams, and architecture design to generate a dependency file requisite for initializing the development environment. This step is followed by the deployment of a standard installation command utilizing the generated file. For Python, the Conda environment manager is employed; for Java and JavaScript, Gradle and NPM are utilized respectively.222 It is noteworthy that our evaluation framework does not contain an Environment Setup task for C/C++ due to the absence of a universally acknowledged and user-friendly dependency management system for these languages (Miranda & Pimentel, 2018). Setting up an environment often encounters challenges such as missing or outdated dependencies, along with version conflicts, all of which must be resolved to ensure a seamless development environment. Our research aims to investigate the potential of LLMs in automating this cumbersome process, thereby enhancing production efficiency.
@@
Figure 2: A typical workflow within DevBench showing building a codebase to develop a TextCNN model for movie review classification.
@@ 3.3 Task 3: Implementation
For this task, models are provided with the PRD, UML diagrams and architecture design, and are then instructed to develop code for each source code file as specified in the architecture design. Diverging from existing benchmarks on repository-level code generation (Liu et al., 2023b), the implementation task in DevBench is dedicated to assessing LLMs in generating an entire code repository from scratch. While similar in spirit to systems like MetaGPT (Hong et al., 2023) and ChatDev (Qian et al., 2023a, b), which derive output from brief requirement descriptions typically under 100 words, our approach is distinct. These systems indeed incorporate software design phases, yet the brevity and lack of detail in their input requirements pose significant challenges in evaluating the extent to which LLMs adhere to these initial specifications. In contrast, real-world software development scenarios often entail more detailed requirements to ensure the final product aligns precisely with expectations, with acceptance testing employed for verification (as discussed in the following subsection).
To more accurately simulate real-world development practices and ensure rigorous evaluation, the implementation task in DevBench involves supplying LLMs with comprehensive inputs including the PRD, UML class and sequence diagrams, and architecture design. The models are then prompted to generate code files. Given the constraint of output length, we adopt a sequential generation approach, prompting the models to produce one code file per interaction. Regarding the planning aspect, recent studies have explored prompting LLMs or training a specific planner (Yao et al., 2023; Besta et al., 2023; Wang et al., ). Considering the structured nature of code files and the inherent dependencies among them, we utilize these dependencies as a clue for effective planning. The generation process is guided to adhere to a partial order derived from a predefined directed acyclic graph, thereby ensuring structured and logical code development. We leave the exploration of planning generated by models themselves for future work.
Evaluation. For the evaluation of the implementation task, an automated testing framework has been developed. This framework, tailored to the specific programming language in use, integrates PyTest for Python, GTest for C++, JUnit for Java, and Jest for JavaScript. The evaluation procedure involves executing reference acceptance and unit tests within a predefined reference environment and the evaluation metric is determined by the pass rate of these tests.
3.4 Task 4: Acceptance Testing
3.5 Task 5: Unit Testing
In this phase, the PRD, UML diagrams, and architecture design are furnished to the models to generate unit tests, facilitating a comprehensive understanding of the software. Unit testing serves as a fundamental approach to safeguard code integrity and operational accuracy. Distinguished from broader acceptance testing, unit testing examines individual code segments for adherence to specified functionalities. The increasing reliance on LLMs for streamlining software development processes underscores the criticality of their adeptness in writing effective unit tests, a competency integral to the software’s reliability and overall robustness.
[…]
@@ 3.5 Task 5: Unit Testing
[…]
Table 2: DevBench Statistics. $\dagger$}: DevBench covers a range of domains including NLP, computer vision, deep learning, algorithm implementation, API applications, Database applications, web service (both frontend and backend), and general tools and utilities. $\ddagger$}: We do not include unit testing for JavaScript as checking functional correctness is not applicable to pure static web pages. Correctness of page rendering and user interaction handling is checked using acceptance tests. $\ast$}: Interfaces with thirty-party libraries that are not used to implement the designated functionalities are not supplied with test cases, resulting in relatively lower overall test coverage.
@@
[…]
4 Dataset
4.1 Dataset Construction
4.2 Dataset Statistics
In Table 2, we present an exhaustive statistical breakdown of our datasets. DevBench contains a collection of 22 curated repositories, spanning across four widely-used programming languages (Python, C/C++, Java, JavaScript) and a diverse range of domains. For detailed repositories statistics, refer to Table 8 in Appendix D. The dataset is characterized by its multi-file structure. The Python repositories in our dataset are relatively straightforward, with each repository comprising approximately two files and an average of 276 lines of code. In contrast, the repositories pertaining to statically-typed programming languages, namely C/C++ and Java, are more complex, featuring an increased count of code files and lines. For JavaScript, our usage of the Vue.js framework necessitates that models adeptly navigate the framework’s templates and development paradigms. Consequently, JavaScript repositories exhibit the highest number of code files and lines, posing a substantial challenge for LLMs. Additionally, we have prepared extensive reference acceptance and unit tests for each repository to facilitate rigorous evaluation of the implementation task.
5 A Baseline System
We introduce a baseline system formulated for our DevBench, building upon the foundations of ChatDev (Qian et al., 2023a, b). ChatDev is a virtual, chat-powered software development system that adheres to the conventional waterfall model. It bifurcates the development process into four primary tasks (dubbed as phrases in ChatDev): design, coding, testing, and documentation. Within this system, multiple LLM agents assume diverse roles such as programmers, reviewers, and testers, pertinent to each phase. ChatDev is characterized by its utilization of a chat chain mechanism, which segments each phase into smaller, atomic tasks. This approach enables context-sensitive, multi-turn dialogues between two distinct roles, facilitating the proposal and validation of solutions for individual tasks.
In contrast to ChatDev, our development of baseline system incorporates several features and enhancements. We have restructured the task design to align closely with the evaluation criteria of DevBench. Specifically, this includes the integration of comprehensive input to the system, exemplified by well-structured PRDs. This integration is crucial in addressing and examining the issue of hallucination in controlled experimental settings. Moreover, our baseline system expands upon the capabilities of ChatDev, supporting a wider range of tasks, including standard Object-Oriented programming designs (UML Class and Sequence diagrams), repository planning (architecture design), environment setup, and acceptance testing. A significant advancement in our baseline system is its compatibility with multiple programming languages and their corresponding runtime environments. This feature is coupled with the provision of comprehensive execution feedback to the system.
@@ 5 A Baseline System
[…]
6 Experiments
6.1 Setup
Base Models
We evaluate three prominent pre-trained model families with different model sizes, including both proprietary and open-source models: OpenAI GPT (GPT, 2023), CodeLlama (Roziere et al., 2023), DeepSeek-Coder (Guo et al., 2024). Specifically, our experiments involve GPT-3.5-Turbo, GPT-4-Turbo from OpenAI GPT [^3] We utilize gpt-3.5-turbo-1106,
gpt-4-0125-preview, respectively.
, CodeLlama-Instruct 7B/13B/34B for CodeLlama, and DeepSeek-Coder-Instruct models 1.3B/6.7B/33B for DeepSeek-Coder. For comprehensive details on these models, we refer readers to the original papers.
Prompting Methods
In our baseline system, we explore three prompting methods: No-Review, Normal-Review, and Execution-Feedback. No-Review represents a basic zero-shot prompting with built-in task prompts. Normal-Review involves a dual-role interaction, where the first role generates a solution and the second role reviews and, where necessary, corrects it. This mode is designed to evaluate the impact of review on model performance, in the absence of external inputs. Execution-Feedback, on the other hand, adds more dynamic interaction to the review process. This feedback includes runtime results, error messages, and performance metrics. Such information could enable the reviewing role to make more informed decisions, potentially leading to more accurate and effective solutions. To save API and computational costs, we only review once for all review-involved prompting methods.
Implementation Details
We utilize LMDeploy for the deployment of CodeLlama and DeepSeek-Coder models. [^4] https://github.com/InternLM/lmdeploy Acknowledging the potential for extensive input context in DevBench tasks, we configure the context length to 32K for these models. For the Software Design task, we set the temperature parameter to 0.2, while for the remaining four tasks, we use a temperature of 0. Other hyperparameters in the experiment are maintained at default settings. All code-related tasks are rigorously evaluated in an isolated sandbox environment, utilizing Docker technology.
Evaluation on Software Design
We follow previous work (Zheng et al., 2023a) to conduct a pairwise comparison to determine which response is better, focusing on the metrics of general principles and faithfulness (see the details in Appendix B). To reduce the expenditure of the OpenAI GPT API and human effort, the scope of our evaluation was confined to a subset of our Benchmark. This process involves 192 pairs across eight repositories, eight models, and three sub-tasks. Regarding the LLM judge, we use GPT-4-Turbo as the judge and GPT-3.5-Turbo as the baseline model. To mitigate the issue of position bias (i.e., preferring response at a certain position regardless of the content), the evaluation was executed in a dual mode, evaluating each pair twice in different orders (384 pairs in total), with inconsistent decisions being considered as a tie. For the human evaluation, we shuffle the order of two responses and annotate each pair thrice to obtain the human majority.
Task Unit Testing
[…]
GPT-4-Turbo
    No-Review w/ Src Code 35.1 34.3 (54.8)
    Normal-Review w/ Src Code 22.6 25.8 (68.7)
CodeLlama-34B-Instruct
    No-Review w/ Src Code 10.7 18.3 (73.2)
    Normal-Review w/ Src Code 12.6 27.0 (72.1)
DeepSeek-Coder-33B-Instruct
    No-Review w/ Src Code 27.2 37.9 (75.8)
    Normal-Review w/ Src Code 22.5 35.5 (71.0)
[…]

### v3

# Prompting Large Language Models to Tackle the Full Software Development Lifecycle: A Case Study

@@ Abstract
Recent advancements in large language models (LLMs) have significantly enhanced their coding capabilities. However, existing benchmarks predominantly focused on simplified or isolated aspects of coding, such as single-file code generation or repository issue debugging, falling short of measuring the full spectrum of challenges raised by real-world programming activities. In this case study, we explore the performance of LLMs across the entire software development lifecycle with DevEval, encompassing stages including software design, environment setup, implementation, acceptance testing, and unit testing. DevEval features four programming languages, multiple domains, high-quality data collection, and carefully designed and verified metrics for each task. Empirical studies show that current LLMs, including GPT-4, fail to solve the challenges presented within DevEval. Our findings offer actionable insights for the future development of LLMs toward real-world programming applications.  [^1] Our data and code are available at https://github.com/open-compass/DevEval.
Prompting Large Language Models to Tackle
the Full Software Development Lifecycle: A Case Study
Bowen Li1thanks: Equal contribution., Wenhan Wu211footnotemark: 1thanks: Work done during internship at Shanghai AI Laboratory., Ziwei Tang311footnotemark: 122footnotemark: 2, Lin Shi411footnotemark: 122footnotemark: 2, John Yang5, Jinyang Li6, Shunyu Yao5, Chen Qian7, Binyuan Hui8, Qicheng Zhang1, Zhiyin Yu1, He Du1, Ping Yang9, Dahua Lin1, Chao Peng9thanks: Corresponding authors., Kai Chen133footnotemark: 3 1Shanghai AI Laboratory, 2Nanjing University, 3BUPT, 4Dartmouth College, 5Princeton University, 6The University of Hong Kong, 7Tsinghua University, 8Alibaba Group, 9ByteDance Correspondence: chenkai@pjlab.org.cn and pengchao.x@bytedance.com
1 Introduction
Figure 1: Our DevEval features multiple stages of software development, including software design, environment setup, implementation, and testing (both acceptance and unit testing).
Task Input Output Environment Evaluation
@@ 1 Introduction
Table 1: Task design in DevEval. $\dagger$}: following our modular evaluation protocol, the input for each task are reference. $\ddagger$}: includes UML class and sequence diagrams. $\ast$}: implementation source code is optional for acceptance testing.
@@
Given its practical value and reasoning challenges, programming has become an important domain to deploy and evaluate large language models (LLMs), leading to popular products like GitHub Copilot and benchmarks like HumanEval (Chen et al., 2021) and APPS (Hendrycks et al., 2021). While these earlier coding tasks focused on generating a single code file or even a single method from simple instructions, recent works such as SWE-bench Jimenez et al. (2023) and RepoBench Liu et al. (2023b) evaluate LLMs on repository-level tasks, which feature longer, more involved NL2Code problems. Still, these benchmarks concentrate on narrow aspects of software development, leaving a gap in comprehensive studies that encompass the full software development lifecycle across its various phases.
@@
To address these shortcomings and fill this gap, we present DevEval, a comprehensive case study that mirrors real-world software development. DevEval generally evaluates models on the task of constructing a multi-file codebase starting from a product requirement document (PRD) of detailed specifications. Subscribing to the traditional Waterfall software development model Royce (1987), DevEval breaks down this process into a diverse set of inter-related development stages, i.e., software design, environment setup, implementation, acceptance and unit testing, as visualized in Figure 1 and Table 1. In contrast to previous works, DevEval is the first to evaluate models’ software design and environment setup capabilities. One significant challenge in this study lies in the scarcity of publicly available repositories that include the full range of software development artifacts, particularly design documents and comprehensive testing programs. To overcome this, we curated a collection of 22 repositories across four programming languages (Python, C/C++, Java, JavaScript), spanning various domains such as machine learning, web services, and command-line utilities. By encompassing the multi-faceted, interconnected steps of software development under a single framework, DevEval provides a holistic view of LLMs’ capabilities for automated software production, moving beyond the conventional focus on code completion.
@@
Through a comprehensive experimental study, our findings indicate that tested models struggle significantly with the challenges presented. GPT-4-Turbo achieves the highest scores amongst all evaluated models, yet it obtains less than 10% on our repository-level implementation task. Other tasks prove relatively more manageable for models; however, even GPT-4-Turbo struggles to attain scores exceeding 40% on the more complicated ones. For instance, models generally fail to generate executable tests, with oracle test scores falling below 40%. Despite this, the generated testing code demonstrates potential in code coverage, achieving as high as 79.4% when it is executable. Furthermore, our investigation into different prompting methods shows that prompts with external information and execution feedback could yield notable and consistent improvements. Importantly, qualitative analysis shows that models demonstrate difficulties in handling Makefile and Gradle, configuring function arguments and employing advanced object-oriented programming techniques. Overall, DevEval introduces a novel challenge for existing LLMs, and our investigation sheds light on fundamental issues, paving the way for future research.
@@ 2 Related Work
Initial code generation datasets focused on self-contained, Python-based completion problems Chen et al. (2021); Hendrycks et al. (2021); Austin et al. (2021), while later works expanded complexity by increasing language coverage Zheng et al. (2023b); Cassano et al. (2023), enhancing execution-based test coverage Liu et al. (2023a); Wang et al. (2022), incorporating dependencies Lai et al. (2022); Ding et al. (2023); Liu et al. (2023c), introducing interactive environments Yin et al. (2022); Yang et al. (2023), and developing short-form task suites Lu et al. (2021); Muennighoff et al. (2023). DevEval  aligns with recent repository-scale coding works Jimenez et al. (2023); Liu et al. (2023b); Zhang et al. (2023). However, our work distinctly evaluates LLMs’ ability to create entire codebases from extensive natural language descriptions, offering a more comprehensive test of LLM capabilities. In automatic programming, role-play frameworks with communicative agents Hong et al. (2023); Li et al. (2023); Qian et al. (2023a, b) like MetaGPT Hong et al. (2023) integrate structured workflows and task modularity, similar to our approach. Compare against MetaGPT, DevEval stands out by providing a structured evaluation across all phases of the software development lifecycle. Additionally, there is a growing trend in agent-based frameworks specially targeting software engineering tasks Yang et al. (2024); Zhang et al. (2024); Liu et al. (2024); Wang et al. (2024a).
3 DevEval
In this section, we discuss the design of DevEval, including task specifications and the evaluation criteria (summarized in Table 1). Adhering to the modular approach, our design utilizes reference inputs for each task. This strategy enables and concentrates on evaluating the efficacy of models in executing specific tasks. 222 Notably, our framework can be adeptly configured to facilitate end-to-end evaluations, utilizing the intermediate outputs generated by models across multiple tasks. Moreover, DevEval  can also function in a Copilot mode, empowering human users to intervene and refine model outputs, thus enhancing the collaborative synergy between human expertise and automated systems.
@@ 3.1 Task 1: Software Design
During this phase, the model is tasked with interpreting the PRD to create the design of the system before development. The first subtask involves the generation of the UML class diagram using Mermaid syntax, which models the structural aspects of software systems, detailing the classes within the system, their attributes, operations and the relationship among them. Class diagrams help developers understand the system’s foundation before development begins. The second subtask focuses on the creation of the UML sequence diagram using Mermaid syntax. These diagrams clarify the collaboration among components of the system by mapping out the interaction between objects and processes over time, illustrating the sequence of messages exchanged between objects to implement the system’s functionality. The final subtask fomulates architectural designs using hierarchical file tree structures, aiming to establish a structured framework for the source code, build scripts and necessary auxiliary files. The architectural design ensures a cohesive structure for coding, testing and maintenance.
Evaluation. Since the Software Design tasks are open-ended, we employ the LLM-as-a-judge approach (Zheng et al., 2023a; Wang et al., 2023; Chiang and Lee, 2023) to conduct the automatic evaluation. The evaluation is anchored by two principal metrics: general principles and faithfulness.
@@ 3.2 Task 2: Environment Setup
In the second phase of DevEval, models are provided with the PRD, UML diagrams, and architecture design to generate a dependency file for initializing the development environment. This step is followed by the deployment of a standard installation command utilizing the generated file. For Python, the Conda environment manager is employed; for Java and JavaScript, Gradle and NPM are utilized respectively.333 It is noteworthy that our evaluation does not contain an Environment Setup task for C/C++ due to the absence of a universally acknowledged and user-friendly dependency management system for these languages Miranda and Pimentel (2018). Setting up an environment often encounters challenges such as missing or outdated dependencies, along with version conflicts, all of which must be resolved to ensure a seamless development environment. Our research aims to investigate the potential of LLMs in automating this cumbersome process, thereby enhancing production efficiency.
@@ 3.3 Task 3: Implementation
For this task, models are provided with the PRD, UML diagrams and architecture design, and are then instructed to develop code for each source code file as specified in the architecture design. Diverging from existing benchmarks on repository-level code generation (Liu et al., 2023b), the implementation task in DevEval  is dedicated to assessing LLMs in generating an entire code repository from scratch. A key innovation in our investigations is the detailed level of requirements provided to the LLMs. In contrast to similar systems like MetaGPT (Hong et al., 2023) and ChatDev (Qian et al., 2023a, b), which generate outputs from brief requirement descriptions typically under 100 words, DevEval offers document-level detail to guide the models. This approach ensures the products are more precisely aligned with expectations, and further acceptance testing is employed for verification. As a result, our evaluations better reflect real-world software development scenarios where detailed requirements are essential to capture complex specifications and ensure product quality.
To more accurately simulate real-world development practices and ensure rigorous evaluation, the implementation task in DevEval  involves supplying LLMs with comprehensive inputs including the PRD, UML class and sequence diagrams, and architecture design. The models are then prompted to generate code files. Given the constraint of output length, we adopt a sequential generation approach, prompting the models to produce one code file per interaction. Regarding the planning aspect, recent studies have explored prompting LLMs or training a specific planner (Yao et al., 2023; Besta et al., 2023; Wang et al., 2024b). Considering the structured nature of code files and the inherent dependencies among them, we utilize these dependencies as a clue for effective planning. The generation process is guided to adhere to a partial order derived from a predefined directed acyclic graph, thereby ensuring structured and logical code development. We leave the exploration of planning generated by models themselves for future work.
Evaluation. For the evaluation of the implementation task, an automated testing framework has been developed. This framework, tailored to the specific programming language in use, integrates PyTest for Python, GTest for C++, JUnit for Java, and Jest for JavaScript. The evaluation procedure involves executing reference acceptance and unit tests within a predefined reference environment. Then the evaluation metric is determined by the pass rate of these tests.
Python C/C++ Java JavaScript
Domain
@@
91.8 95.0 64.9 -
Table 2: DevEval Statistics. $\dagger$}: DevEval covers a range of domains including NLP, computer vision, deep learning, algorithm implementation, API applications, Database applications, web service (both frontend and backend), and general tools and utilities. $\ddagger$}: We do not include unit testing for JavaScript as checking functional correctness is not applicable to pure static web pages. Correctness of page rendering and user interaction handling is checked using acceptance tests. $\ast$}: Interfaces with thirty-party libraries that are not used to implement the designated functionalities are not supplied with test cases, resulting in relatively lower overall test coverage.
3.4 Task 4: Acceptance Testing
3.5 Task 5: Unit Testing
In this phase, models are given the PRD, UML diagrams, and architecture design to generate unit test codes that facilitate a comprehensive understanding of the software. Unit testing serves as a fundamental approach to safeguard code integrity and operational accuracy. Distinguished from broader acceptance testing, which focuses on the overall functionality and viability of the software, unit testing examines individual code segments for adherence to specified functionalities. The increasing reliance on LLMs for streamlining software development processes underscores the criticality of their adeptness in writing effective unit tests, a competency integral to the software’s reliability and overall robustness.
@@ 3.5 Task 5: Unit Testing
3.6 Dataset
The dataset construction process involved three phases: repository preparation, code cleanup, and document preparation. We first selected high-quality repositories from a GitHub dump, applying filters to ensure manageable complexity for evaluation. Postgraduate student annotators then set up the environments, executed the code to verify functionality, and cleaned the repositories by removing unnecessary files and running or creating unit and acceptance tests to ensure oracle test standards. Finally, they prepared standard software design documents, including UML diagrams and architecture designs, following specific guidelines. The curated dataset consists of 22 repositories across Python, C/C++, Java, and JavaScript, with varying complexities and multi-file structures, as detailed in Table 2. Appendix B provides more details.
4 Experiments
4.1 Setup
Models and the Baseline System
We evaluate three prominent pre-trained model families with different model sizes, including both proprietary and open-source models: OpenAI GPT (OpenAI, 2023), CodeLlama (Roziere et al., 2023), DeepSeek-Coder (Guo et al., 2024). Specifically, our experiments involve GPT-3.5-Turbo, GPT-4-Turbo from OpenAI GPT [^4] We utilize gpt-3.5-turbo-1106,
gpt-4-0125-preview, respectively.
, CodeLlama-Instruct 7B/13B/34B, and DeepSeek-Coder-Instruct models 1.3B/6.7B/33B. Regarding the baseline system, we extend ChatDev Qian et al. (2023a, b) for DevEval, adding support for UML diagrams, architecture design, environment setup, and multi-language execution, with structured PRDs and feedback to reduce hallucinations. Further details are provided in the Appendix C.
@@ 4.1 Setup
Table 3: Results of different prompting methods of the Implementation task on a subset of DevEval. $\P$: all results are averaged across all repositories and weighted by the number of code lines, which measures the difficulty of each repository.
Prompting Methods
In our baseline system, we explore three prompting methods: No-Review, Normal-Review, and Execution-Feedback. No-Review represents a basic zero-shot prompting with built-in task prompts. Normal-Review involves a dual-role interaction, where the first role generates a solution and the second role reviews and, where necessary, corrects it. This mode is designed to evaluate the impact of review on model performance, in the absence of external inputs. Execution-Feedback, on the other hand, adds more dynamic interaction to the review process. This feedback includes runtime results, error messages, and performance metrics. Such information could enable the reviewing role to make more informed decisions, potentially leading to more accurate and effective solutions. For computational efficiency, we conduct a single review for all review-involved prompting methods.
Task Unit Testing
Evaluation Metric (%) Oracle Test§ Coverage\$
GPT-4-Turbo
      No-Review w/ Src Code 35.1 34.3 (54.8)
      Normal-Review w/ Src Code 22.6 25.8 (68.7)
CodeLlama-34B-Instruct
      No-Review w/ Src Code 10.7 18.3 (73.2)
      Normal-Review w/ Src Code 12.6 27.0 (72.1)
DeepSeek-Coder-33B-Instruct
      No-Review w/ Src Code 27.2 37.9 (75.8)
      Normal-Review w/ Src Code 22.5 35.5 (71.0)
Table 4: Results of different prompting methods of the Unit Testing task on a subset of DevEval. $\S$§: the Oracle Test results are averaged across all repositories and weighted uniformly. \$currency-dollar\\$\$: the results on the left side are averaged across all repositories and weighted uniformly, showing the overall scores. The results on the right side in the parenthesis are averaged across all valid repositories and weighted uniformly, where models have generated executable testing code.
@@
Table 5: Results of different prompting methods of the Acceptance Testing task on a subset of DevEval. $\S$§: all results are averaged across all repositories and weighted uniformly.
4.2 Main Results
Results across prompting methods
We first conduct experiments to examine the effects of different prompting methods on a subset of DevEval using three representative models: GPT-4-Turbo, CodeLlama-34B-Instruct, and DeepSeek-Coder-33B-Instruct.